OpenStudy (dmezzullo):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (anonymous):

Distribute. −3(x + 3) = −3(x + 1) − 5 -3x - 9 = -3x - 3 - 5 Combine like terms. -3x - 9 = -3x - 3 - 5 -3x - 9 = -3x - 8 0 = 1 So no real solutions.

OpenStudy (dmezzullo):

tyvm medal?

OpenStudy (anonymous):

lol welcome (: